Printing press model A can print 70 books per day and model B can print 55 per day. If there are 14 total presses and together they print 905 books per day, how many books does each model print per day?
Answer by ankor@dixie-net.com(22740)

Printing press model A can print 70 books per day and model B can print 55 per day.
If there are 14 total presses and together they print 905 books per day, how many books does each model print per day?
:
No. of presses equation
A + B = 14
B = (14-A); use this form for substitution
:
Output equation
70A + 55B = 905
Replace B with (14-A)
70A + 55(14-A) = 905
70A + 770 - 55A = 905
70A - 55A = 905 - 770
15A = 135
A = 135/15
A = 9 presses
:
You can find the number of B presses, Check the solutions in the output equation